What is the magnitude of the electric field at a field point 2m from a point charge of 4nC


1
Expert's answer
2022-01-25T06:49:49-0500

Given:

"r=2\:\rm m"

"q=4\:\rm nC=4*10^{-9}\: C"

"k=9*10^{9}\: \rm N\cdot m^2/C^2"


The electric field of a point charge is given by

"E=k\frac{q}{r^2}=9*10^9*\frac{4*10^{-9}}{2^2}=9\:\rm N/C"